OpenStudy (thatonegirl_):

How to solve e^-x=5?

OpenStudy (anonymous):

Are you familiar with logarithms?

OpenStudy (thatonegirl_):

sorta, mostly natural log and base 10.. i forget e. @swagmaster47

OpenStudy (anonymous):

You can take the log of both sides so you have\[-xlog(e=\log(5\]

OpenStudy (anonymous):

Hi there! I am going to give a similar example. \[e^{2x}=2\]\[\ln(~e^{2x}~)=\ln(2)\]\[2x~\ln(~e)=\ln(2)\]\[2x~(1)=\ln(2)\]\[2x=\ln(2)\]\[x=2/\ln(2)\]

OpenStudy (isaiah.feynman):

The North Korean president is good in maths!

OpenStudy (anonymous):

IF, lol:)

OpenStudy (isaiah.feynman):